OpenStudy (anonymous):

the perimeter of a rectangle is 102 inches. the length is 5 inches greater that the width. what is an equation for this sentence.

OpenStudy (anonymous):

let the width be x length is x + 5 (x + x + 5)*2 = 102 2(2x+5) = 102 2x + 5 = 51 2x = 46 x = 23

OpenStudy (anonymous):

thank you